Flamo:

The legs of a right triangle have lengths of 4 and 5 units. Find the length of the hypotenuse.

Flamo:

Okay so.

Flamo:

a^2 + b^2 = c^2

Angle:



Flamo:

The legs are 4 and 5

Angle:

yup

Flamo:

Ye

Flamo:

So, 4^2 + 5^2 = c^2

Flamo:

16 + 25 = c^2 =

Flamo:

41 = c^2

Flamo:

So its the Square root of 41. :D

Angle:

perfect! you did great!

Flamo:

Ye!

Flamo:

Ty!